theyadoreshayyy:

In △JKL, solve for x. 66.73 74.89 15.44 38.16

theyadoreshayyy:

snowflake0531:

Do you know soh cah toa?

theyadoreshayyy:

no

snowflake0531:

\[\sin(\theta) = \frac{opposite}{hypotenuse} ~~~~~ \cos(\theta)= \frac {adjacent}{hypotenuse} ~~~~~ \tan(\theta) = \frac {opposite}{adjacent} ~ the~adjacent~and~opposite~angles~are~coming~from~the~reference~\angle~while~the~hypotenuse~is~just~the~regular~hypotenuse\]

snowflake0531:

oops o-O

smith123:

ima cracker

snowflake0531:

So look from angle J, the 27, which of the 2 do you have out of adjacent, opposite and hypotenuse

theyadoreshayyy:

34, and x

snowflake0531:

but like what are they adjacent opposite hypotenuse...?

theyadoreshayyy:

the sine one

snowflake0531:

Yes so can you write out the equation using sin(theta) = opposite/hypotenuse ?

theyadoreshayyy:

<X=sinx= opposite over hypotenuse = x over 34

theyadoreshayyy:

i think thats right, im not sure

snowflake0531:

\[\sin27 = \frac {34}{x}\]

theyadoreshayyy:

so i was right ?

snowflake0531:

Well, now find x

theyadoreshayyy:

how would i find it ?

snowflake0531:

1 attachment
theyadoreshayyy:

thank yuuu

snowflake0531:

yw~